OpenStudy (anonymous):

>>>DEATH STAR Question: Find the smallest a and b when a and b = positive integer QUESTION:VVVVVVVV

OpenStudy (anonymous):

\[\frac{a}{b}=\frac{1}{1^{4}+1^{2}+1}+\frac{2}{2^{4}+2^{2}+1}+...+\frac{100}{100^{4}+100^{2}+1}\]

OpenStudy (anonymous):

have fun :)

OpenStudy (raden):

use this :  THUS = 1/2 * (100^2+100)/(100^2+100+1) = 1/2 * 10,100/10,101 = 5,000/10,101